Grade 12General Physics

kinetic energy of a satellite is x then what is the time period proportional to
in my package it says its proportional to x^-3/2
how?

SHREYA GAUTAM
10 Years agoGrade 12

1 Answer

Vikas TU
10 Years ago
From closest distance approach,
K.E = P.E. = Gm1m2/r = x (given)
from here,
we conclude that 
x is inversely prop. to r or vice versa.
r = k/x …...................(1)
Now from, kepler’s  time period law,
T^2 is directly proprotiional to the r^3
or
r is directly proportional to T^(2/3)
.i.e
r = kT^(2/3) …..................(2)
from 1 and 2 eqn.
T is inversely proportional to x^3/2
or 
 
proportional to x^-3/2
Hence proved.